OpenStudy (anonymous):

Short answer. Find the product -4x(-7x^3+4x^2+8x-6) can someone help me please

OpenStudy (anonymous):

when you multiply ax * bx =(a*b)x^2 for example: (-4x)*(-7x^3)=28x^4

OpenStudy (anonymous):

general formula:\[ax ^{n}*bx ^{n}=(a*b)x ^{n+m}\]

OpenStudy (anonymous):

do i multiply everything in the parentheses by -4x

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

i still am a little confused with this with all the powers

OpenStudy (anonymous):

please work this one out step by step so i understand better

jimthompson5910 (jim_thompson5910):

When you multiply something like -4x and -7x^3, you multiply the coefficients -4 and -7 to get (-4)*(-7) = 28 and you add the exponents: 1+3 = 4 So this means that -4x times -7x^3 is 28x^4 Do the rest for the remaining terms.

jimthompson5910 (jim_thompson5910):

Note: -4x is the same as -4x^1

OpenStudy (anonymous):

ok well thank you very much

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

i totally understand now. cool

jimthompson5910 (jim_thompson5910):

that's great, I'm glad you do

OpenStudy (anonymous):

i really like this open study. It has really helped me alot. Everyone has been so helpful to me!!!!!!

jimthompson5910 (jim_thompson5910):

yes it's really a great site

OpenStudy (anonymous):

how do you get a pic on here

jimthompson5910 (jim_thompson5910):

Click "settings" up at the top

jimthompson5910 (jim_thompson5910):

then click on the link "change picture"